# ReCLIP++: Learn to Rectify the Bias of CLIP for Unsupervised Semantic Segmentation

### v1

@@ III-A Baseline: Directly Segment with CLIP
Following MaskCLIP [23], we adapt pre-trained CLIP [2] (ViT-B/16) to the semantic segmentation task. We remove the query and key embedding layers of last attention but reformulate the value embedding layer and the last linear projection layer into two respective $1\times 1$ convolutional layers. Therefore, the image encoder can not only generate local feature for dense prediction, but also keep the visual-language association in CLIP by freezing its pre-trained weights. We forward image $I$ through the image encoder and obtain patch-level feature $Z\in\mathbb{R}^{n\times D}$ ($n$ is the number of patches).
@@
Each text $Q_{i}$ in Query $Q=\{Q_{1},Q_{2},\cdots,Q_{C}\}$ ($C$ is the number of classes) is an ensemble of several manually designed templates, e.g., “a good/large/bad photo of a $[CLS]$” where $[CLS]$ denotes a specific class name. Passing $Q_{i}$ through the text encoder, we obtain different text embeddings for different manually designed templates. We average all those text embeddings as the embedding of $Q_{i}$. Then the embeddings of all the $Q_{i}$ in $Q$ make up the query text feature $W_{q}\in\mathbb{R}^{C\times D}$. We treat text feature $W_{q}$ as the weight of segmentation head to perform 1 $\times$ 1 convolution. By sending feature $Z$ to the segmentation head, we get a Query logit map $M_{q}\in\mathbb{R}^{n\times C}$. Then the segmentation mask can be predicted by $\arg\max$ operation on $M_{q}$ across class channels.
III-B Bias Extraction
@@
where each $[v_{l}](l\in\{1,...,L\})$ is a vector with dimension $D$ and serves as a word embedding. The $L$ is a constant representing the number of word embeddings to learn. Totally, there are $77$ word embeddings for a text input of CLIP. Among these, two word embeddings are used for class name and two for indicating start and end of a text. Thus, we set $L$ to $73$.
@@
Space-preference Bias Encoding. In ViT [65], positional embedding (PE) is important for encoding spatial information into feature. Thus, we assume the space-preference bias should depend on the positional embedding (PE) and we choose to learn a projection of PE to encode the space-preference bias. The projection network of PE is designed as a 1-layer $1\times 1$ convolutional network. Then we project PE $p$ by the designed convolutional network to obtain positional feature $W_{p}\in\mathbb{R}^{n\times D}$. During the training process, the projection network is optimized to encode the space-preference bias with respect to each patch of image. Note that the encoding process of space-preference bias is independent to that of class-preference bias because the encoding of space-preference bias doesn’t rely on any class information.
@@
Finally, $M_{b}$ represents the bias that we aim to remove from the predicted logit map $M_{q}$ of original CLIP.
@@
Differently, in ReCLIP++, we first encode class-preference bias and space-preference bias into Reference feature and positional feature respectively. Reference feature only depends on class-related information (i.e., learnable Reference text), while positional feature only depends on space-related information (i.e., PE). Thus, in this way, we may independently encode two kinds of biases and largely avoid the interference. As a result, with ReCLIP++, we realize more effective bias rectification and achieve better segmentation results (see the Sec. IV for more details).
III-C Rectified Mask Generation
By keeping Query prompt $Q$ fixed, the Query logit map $M_{q}$ represents the natural prediction ability of CLIP model, which may contain class-preference bias and space-preference bias. With bias extraction module, we obtain $M_{b}$ which represents the bias that we aim to remove from CLIP predictions. In this paper, we perform a simple element-wise subtraction between $M_{q}$ and $M_{b}$ to generate the rectified logit map $M$, i.e.,
@@ III-C Rectified Mask Generation
From a high-level, this operation can be interpreted as subtracting the “bias” from the predictions with Query text feature $W_{q}$ as segmentation head.
@@
To generate the semantic mask $\hat{M}\in\mathbb{R}^{n\times C}$ for each category, a natural way is to employ $\arg\max$ operation to $M_{o}$ across class channels and generate the one-hot label for each spatial location. Then, each channel of $\hat{M}$ corresponds to the mask for a specific category. However, as the $\arg\max$ operation is not differentiable, we are not able to optimize $\hat{M}$ end-to-end through gradient back-propagation during the training process. Thus we utilize a Gumbel-Softmax trick [66] to generate the semantic mask for each category with $\tau_{1}$ as the temperature, i.e.,
@@
In this way, with our designed supervision (Sec. III-D), the gradient can be back-propagated through $\hat{M}$ to encourage more effective bias rectification and better semantic mask generation.
III-D Bias Rectification via Contrastive Loss
@@
Image-level multi-label hypothesis generation. In order to compute the contrastive loss, we need to know what classes exist in a specific image. However, this information is not available in unsupervised semantic segmentation setting. In this section, we propose to generate image-level multi-label hypothesis which means the set of classes that potentially exist in an image, to facilitate the following contrastive loss calculation.
In ReCLIP [27], we directly calculate the similarity scores between the class-level text features of manually designed prompts, e.g., “a photo of a $[CLS]$”, and the image-level visual feature extracted by CLIP. We select the set of classes whose scores are higher than a threshold as the hypothesis. As CLIP is pre-trained with aligning image-level visual feature with text feature, it may focus on the most salient objects in an image and fail to recognize every existing object, rendering the multi-label hypothesis generated by ReCLIP less accurate.
In ReCLIP++, we design a new strategy to generate more accurate multi-label hypothesis for an image. Given an image, we firstly split it into overlapped crops. Specifically, we slide a window across an image with a certain stride. In our implementation, we set the window width and height to $r$ times the width and height of the input image and set the horizontal and vertical stride to half of the window width and height respectively. For each crop, we forward it through the visual encoder of CLIP to obtain its global visual feature, and compute the similarity scores between the global visual feature and the Query feature $W_{q}$. We treat the class with the highest score as the detected class for this crop. Then for a given image, among all of its crops, we calculate the frequency $f(k)$ of the $k$-th class, $k\in\{1,2,...,C\}$, being detected. We choose the class whose frequency is higher than the threshold $t$ as the multi-label hypothesis $\mathcal{H}$ for an image, i.e.,
@@
Bias Rectification via Contrastive Loss We then design a contrastive loss to supervise the rectification process via aligning masked features with text features of corresponding classes. Specifically, we apply $\hat{M}$ to the feature $Z$ and perform a global average pooling to get the class-level masked features $Z_{g}\in\mathbb{R}^{C\times D}$, which encode the features of regions with different classes. We compute the similarities between $Z_{g}$ and text features of all the categories. As the masked features $Z_{g}$ represent the objects of interest without context, we utilize $W_{q}$ generated by text input $Q$ for similarity computation. With generated $\mathcal{H}$, the contrastive loss is computed as
@@
A better modeling of bias yields more accurate estimations of object masks. Then the masked features of objects are more aligned with corresponding text features. As a result, the contrastive loss (Eq. (7) will therefore be lower. In contrast, a worse modeling of bias results in higher contrastive loss. Thus, minimizing Eq. (7) will drive the model to update towards making more accurate mask predictions, i.e., rectifying the bias of CLIP when adapting CLIP to the downstream USS task.
@@
Inference. After the rectification stage, we obtain a rectified CLIP. For the inference with rectified CLIP, we model both biases in bias logit map $M_{b}$ by the Bias Extraction Module as III-B and obtain the rectified output $M_{o}$ by the Rectified Mask Generation Module as III-C. Finally, we simply employ $\arg\max$ operation to $M_{o}$ across class channels to generate the rectified masks as final predictions.
III-E Distillation for Enhanced Results
Since CLIP is not specifically designed for segmentation tasks, we choose to distill the knowledge from the rectified CLIP (teacher) to the advanced segmentation architecture (student) to enhance the feature representations and finally improve the segmentation performance. In our paper, we choose DeepLab V2 [67] as the student network which directly inherits and freezes the Query segmentation head, bias logit map and the Rectified Mask Generation Module from the rectification stage. In detail, for the student, the visual feature $\hat{Z}$ is generated by the student backbone. Passing the visual feature through the inherited Query segmentation head, we obtain the Query logit map. We input the Query logit map and the inherited bias logit map into the inherited Rectified Mask Generation Module to generate the rectified predictions $M^{D}$ of student. In our design, the rectified CLIP works as a teacher to guide the feature learning of DeepLab V2 with three designed loss terms.
Mask-guided loss. We directly exploit segmentation masks generated from the rectification stage as pseudo labels. We compute a cross-entropy loss between the pseudo labels $\tilde{M}$ and predictions $M^{D}$ from DeepLab V2, i.e.,
$\mathcal{L}_{mask}=-\frac{1}{HW}\sum^{H}_{i}\sum^{W}_{j}\log P_{ij}(\tilde{M}_{ij})$ (8)
where $H$ and $W$ denote the height and width of an image. The $P_{ij}(\tilde{M}_{ij})=\text{Softmax}(M^{D}_{ij})_{\tilde{M}_{ij}}$, which means the predicted probability with respect to the pseudo label $\tilde{M}_{ij}$.
Feature-guided loss. After obtaining feature $Z$ from visual backbone of rectified CLIP and $\hat{Z}$ from student visual backbone, we resize two features to the same shape of original image by bilinear interpolation. We then perform a L1 loss between two features
$\mathcal{L}_{feat}=||Z-\hat{Z}||_{1}$ (9)
Therefore, visual feature from student network is aligned with visual feature of rectified CLIP. Benefiting from the alignment between textual and visual features of CLIP, this loss provides an important basis for the following text-guided loss.
Text-guided loss. We adopt the same strategy as the Rectification stage (Sec. III-D) to compute contrastive loss as our text-guided loss $\mathcal{L}_{text}$. The only difference is that we use feature $\hat{Z}$ and masks generated by the student network to obtain the corresponding class-level masked features.
Total distillation loss. Total distillation loss $\mathcal{L}_{distill}$ is calculated as follows, where $\alpha$ and $\beta$ are both constants. Effect of each loss term is studied in Sec. IV-D.
$\mathcal{L}_{distill}=\mathcal{L}_{mask}+\alpha\mathcal{L}_{feat}+\beta\mathcal{L}_{text}.$ (10)
IV Experiments
IV-A Setup
@@
Datasets. We conduct experiments on five standard benchmarks for semantic segmentation, including PASCAL VOC 2012 [1], PASCAL Context [3], ADE20K [28], Cityscapes [29] and COCO Stuff [30]. PASCAL VOC 2012 (1,464/1,449 train/validation) contains 20 object classes, while PASCAL Context (4,998/5,105 train/validation) is an extension of PASCAL VOC 2010 and we consider 59 most common classes in our experiments. ADE20K (20,210/2,000 train/validation) is a segmentation dataset with various scenes and 150 most common categories are considered. Cityscapes (2,975/500 train/validation) consists of various urban scene images of 19 categories from 50 different cities. COCO Stuff (118,287/5,000 train/validation) has 171 low-level thing and stuff categories excluding background class. Following previous evaluation protocol of unsupervised semantic segmentation methods [24, 45], we use 27 mid-level categories for training and inference.
@@
Implementation details. For the image encoder of CLIP, we adopt ViT-B/16 as our visual backbone. For the text encoder of CLIP, we adopt Transformer [68]. During the whole training period, we keep both of the encoders frozen. We use conventional data augmentations including random cropping and random flipping. For both stages, we use a SGD [69] optimizer with a learning rate of 0.01 and a weight decay of 0.0005. We adopt the poly strategy with the power of 0.9 as the learning rate schedule. We set $r=1/6$ for all the datasets. We set $t=7\%$ for VOC and Cityscapes, and $t=0$ for the rest datasets. For distillation, we set $\alpha=\beta=0.5$. In our experiment, we report the mean intersection over union (mIoU) as evaluation metric. More details for training can be found in our appendix.
Baselines. We compare with previous USS methods to verify the superiority of our method, including MaskCLIP(+) [23], CLIP-S4 [26], ReCo [24], EAGLE [70], DepthSeg [71], etc. From the technical view, although our method falls into the category of text-guided segmentation, we choose not to directly compare with text-guided methods which are designed for zero-shot or open-vocabulary setting (e.g., GroupViT [57], TCL [63] and ViewCo [61]). Two main differences exist between the zero-shot/open-vocabulary setting and the unsupervised setting: 1) in zero-shot/open-vocabulary, the category sets for training and inference are non-overlapped, but in USS the category sets are shared; 2) text-guided zero-shot/open-vocabulary methods usually rely on large-scale image-text pairs or class-agnostic masks to supervise the training, but in USS we only utilize the images without any types of annotations to train the model.
IV-B Comparison with previous state-of-the-arts
TABLE I: Comparison with non-language-guided USS methods on various benchmarks.
Model VOC Cityscapes COCO Stuff
IIC [15] 9.8 6.4 6.7
SegSort [22] 11.7 / 16.4
DSM [49] 37.2 / /
HSG [17] 41.9 / 23.8
ACSeg [19] 47.1 / 16.4
MaskContrast [20] 49.6 / /
EAGLE [70] / 22.1 27.2
DepthSeg [71] / 23.1 29.0
ReCLIP (LP) 76.1 49.3 26.2
ReCLIP++ (rec.) 73.5 21.7 21.2
ReCLIP++ 79.2 26.9 23.2
ReCLIP++ (LP) 80.3 54.1 29.5
We compare our method with both previous non-language-guided USS methods (Table I) and CLIP-based USS methods (Table II). In all tables, we use “ReCLIP++ (rec.)” to denote the segmentation results after the rectification stage and “ReCLIP++” to denote the segmentation results after rectification and distillation. Results from previous methods are directly cited from the original papers, except that † means the results are obtained by running the officially-released source code. All the numbers reported are expressed as percentages.
As shown in Table I, ReCLIP++ shows remarkably better segmentation results compared with conventional non-language-guided USS methods. These non-language-guided USS methods cannot directly attach a label to the discovered segments and thus rely on some post-processing, e.g., linear probing, to evaluate the segmentation performance. Instead, ReCLIP++ can directly assign a class label to each pixel under the guidance of language. In order to further evaluate the strength of features extracted from ReCLIP++ after both rectification and distillation stage, we also report our results of linear probing (LP). Following MaskContrast [20], we fix our model to generate pixel-level embedding and train a linear classifier to generate semantic segmentation masks based on ground-truth annotations. The results of LP also prove that ReCLIP++ yields better features and achieves obvious gains.
In Table II, we also make a comparison with typical CLIP-based methods, including MaskCLIP, CLIPpy, ReCo and CLIP-S4. These methods all consider the same CLIP-based USS setting with ReCLIP++. The results show ReCLIP++ performs favorably against previous CLIP-based methods. For example, ReCLIP++ (rec.) outperforms MaskCLIP by 24.0%, 12.3%, 4.0%, 1.9% and 7.3% respectively on the five datasets, while after distillation ReCLIP++ outperforms MaskCLIP+ by 9.2%, 3.4%, 2.3%, 1.7% and 3.7%. ReCLIP++ also shows better results than CLIPpy by 24.6%, 1.0% and 4.6% on PASCAL VOC, ADE20K and Cityscapes, and outperforms CLIP-S4 by 7.2% and 0.9% on PASCAL VOC and PASCAL Context respectively. Since ReCo employs a “context elimination” (CE) trick which introduces prior knowledge to assist the training, we also report the results of ReCo by removing this trick (ReCo w/o. CE in Table II) for fair comparison. ReCLIP++ performs favourably against ReCo, e.g., on PASCAL VOC and PASCAL Context, ReCLIP++ outperforms ReCo by 24.0%, 8.3% respectively; on COCO Stuff, ReCLIP++ outpferforms ReCo (w/o. CE) by 2.4%.
TABLE II: Comparison with CLIP-based unsuperivsed semantic segmentation methods on various benchmarks. † means the results are obtained by running the officially released source code.
Method VOC Context ADE Cityscapes Stuff
MaskCLIP [23] $49.5^{\dagger}$ 21.7 $9.5^{\dagger}$ $19.8^{\dagger}$ $13.6^{{\dagger}}$
MaskCLIP+ [23] $70.0^{{\dagger}}$ 31.1 $12.2^{{\dagger}}$ $25.2^{{\dagger}}$ $19.5^{{\dagger}}$
CLIPpy [25] 54.6 / 13.5 22.3 /
ReCo [24] $55.2^{{\dagger}}$ 26.2 / 19.3 26.3
ReCo (w/o. CE) [24] $54.8^{{\dagger}}$ $23.1^{{\dagger}}$ / $18.2^{{\dagger}}$ $20.8^{{\dagger}}$
CLIP-S4 [26] 72.0 33.6 / / /
ReCLIP (rec.) [27] 58.5 25.8 11.1 16.2 14.6
ReCLIP [27] 75.8 33.8 14.3 19.9 20.3
ReCLIP++ (rec.) 73.5 34.0 13.5 21.7 20.9
ReCLIP++ 79.2 34.5 14.5 26.9 23.2
IV-C Comparison with Conference Version
TABLE III: Comparison with ReCLIP. The first line is the result of ReCLIP after the rectification stage and following lines are ablations on the technical improvements. Results show that each newly added technical improvement of ReCLIP++ contributes to better results than ReCLIP.
Decoder Label Hypothesis Independent VOC Context
58.5 25.8

### v2

@@ 3.1 Baseline: Directly Segment with CLIP
Following MaskCLIP [20], we adapt pre-trained CLIP [19] (ViT-B/16) to the semantic segmentation task. We remove the query and key embedding layers of last attention but reformulate the value embedding layer and the last linear projection layer into two respective $1\times 1$ convolutional layers. Therefore, the image encoder can not only generate the patch-level visual feature for dense prediction but also keep the visual-language association in CLIP by freezing its pre-trained weights. We forward image $I$ through the image encoder and obtain patch-level feature $Z\in\mathbb{R}^{n\times D}$ ($n$ is the number of patches).
@@
Each text $Q_{i}$ in Query $Q=\{Q_{1},Q_{2},\cdots,Q_{C}\}$ ($C$ is the number of classes) is an ensembling of several manually designed templates, e.g., “a good/large/bad photo of a $[CLS]$”, where $[CLS]$ denotes a specific class name. Passing $Q_{i}$ through the text encoder, we obtain different text embeddings for different manually designed templates. We average all those text embeddings as the embedding of $Q_{i}$. Then the embeddings of all the $Q_{i}$ in $Q$ make up the query text feature $W_{q}\in\mathbb{R}^{C\times D}$. We treat text feature $W_{q}$ as the weight of the segmentation head to perform 1 $\times$ 1 convolution. By sending feature $Z$ to the segmentation head, we get a Query logit map $M_{q}\in\mathbb{R}^{n\times C}$. Then the segmentation mask can be predicted by $\arg\max$ operation on $M_{q}$ across the class channel $C$.
3.2 Bias Extraction
@@
where each $[v_{l}](l\in\{1,...,L\})$ is a vector with dimension $D$ and serves as a word embedding. The $L$ is a constant representing the number of word embeddings to learn. Totally, there are $77$ word embeddings for a textual input of CLIP. Among these, two word embeddings are used for representing class names and two for indicating the start and end of a textual input. Thus, we set $L$ to $73$.
@@
Space-preference Bias Encoding. In ViT [66], positional embedding (PE) is important for encoding spatial information into features. Thus, we assume the space-preference bias should depend on the positional embedding (PE) and we choose to learn a projection of PE to encode the space-preference bias. The projection network of PE is designed as a 1-layer $1\times 1$ convolutional network. Then we project PE $p$ by the designed convolutional network to obtain positional feature $W_{p}\in\mathbb{R}^{n\times D}$. During the training process, the projection network is optimized to encode the space-preference bias with respect to each patch of the image. Note that the encoding process of space-preference bias is independent of that of class-preference bias because the encoding of space-preference bias doesn’t rely on any class information.
@@
Finally, $M_{b}$ represents the final bias that we aim to remove from the predicted query logit map $M_{q}$ of the original CLIP.
@@
Differently, in ReCLIP++, we first encode class-preference bias and space-preference bias into the Reference feature and positional feature respectively. The reference feature only depends on class-related information (i.e., learnable Reference text), while the positional feature only depends on space-related information (i.e., PE). Thus, in this way, we may independently encode two kinds of biases and largely avoid the interference. As a result, with ReCLIP++, we realize more effective bias rectification and achieve better segmentation results (see Sec. 4 for more details).
3.3 Rectified Mask Generation
By keeping Query prompt $Q$ fixed, the Query logit map $M_{q}$ represents the natural prediction ability of the original CLIP model, which may contain class-preference bias and space-preference bias. With the bias extraction module, we obtain $M_{b}$ which represents the bias that we aim to remove from CLIP predictions. In this paper, we perform a simple element-wise subtraction between $M_{q}$ and $M_{b}$ to generate the rectified logit map $M$, i.e.,
@@ 3.3 Rectified Mask Generation
From a high level, this operation can be interpreted as subtracting the “bias” from the predictions with Query text feature $W_{q}$ serving as the weight of the segmentation head.
@@
To generate the semantic mask $\hat{M}\in\mathbb{R}^{n\times C}$ for each category, a natural way is to employ $\arg\max$ operation to $M_{o}$ across the class channel $C$ and generate the one-hot label for each spatial location. Then, each channel of $\hat{M}$ corresponds to the mask for a specific category. However, as the $\arg\max$ operation is not differentiable, we are not able to optimize $\hat{M}$ end-to-end through gradient back-propagation during the training process. Thus we utilize a Gumbel-Softmax trick [67] to generate the semantic mask for each category with $\tau_{1}$ as the temperature, i.e.,
@@
In this way, with our designed supervision (Sec. 3.4), the gradient can be back-propagated through $\hat{M}$ to encourage a more effective bias rectification and a better semantic mask generation.
3.4 Bias Rectification via Contrastive Loss
@@
Image-level multi-label hypothesis generation. In order to compute the contrastive loss, we need to know what classes exist in a specific image. However, this information is not available under the USS setting. In this section, we propose to generate an image-level multi-label hypothesis which means the set of classes that potentially exist in an image, to facilitate the following contrastive loss calculation.
In ReCLIP [24], we directly calculate the similarity scores between the class-level text features of manually designed prompts, e.g., “a photo of a $[CLS]$”, and the image-level visual feature extracted by CLIP. We select the set of classes whose scores are higher than a threshold as the hypothesis. As CLIP is pre-trained to align image-level visual features with text features, it may focus on the most salient objects in an image and fail to recognize every existing object, rendering the multi-label hypothesis generated for ReCLIP less accurate.
In ReCLIP++, we design a new strategy to generate a more accurate multi-label hypothesis for an image. Given an image, we first split it into overlapped crops. Specifically, we slide a window across an image with a certain stride. In our implementation, we set the window width and height to $r$ times the width and height of the input image and set the horizontal and vertical stride to half of the window width and height respectively. For each crop, we forward it through the visual encoder of CLIP to obtain its global visual feature and compute the similarity scores between the global visual feature and the Query feature $W_{q}$. We treat the class with the highest score as the detected class for this crop. Then for a given image, among all of its crops, we calculate the frequency $f(k)$ of the $k$-th class, $k\in\{1,2,...,C\}$, being detected. We choose the classes with frequency higher than the threshold $t$ as the multi-label hypothesis $\mathcal{H}$ for an image, i.e.,
@@
Bias Rectification via Contrastive Loss. We then design a contrastive loss to supervise the rectification process via aligning masked features with text features of corresponding classes. Specifically, we apply $\hat{M}$ to the feature $Z$ and perform a global average pooling to get the class-level masked features $Z_{g}\in\mathbb{R}^{C\times D}$, which encode the features of regions belonging to different classes. We compute the similarities between $Z_{g}$ and text features of all the categories. As the masked features $Z_{g}$ represent the objects of interest without context, we utilize $W_{q}$ generated by text input $Q$ for similarity computation. With generated $\mathcal{H}$, the contrastive loss is computed as
@@
A better modeling of bias yields more accurate estimations of object masks. Then the masked features of objects are more aligned with the corresponding text features. As a result, the contrastive loss (Eq. (7) will therefore be lower. In contrast, a worse modeling of bias results in higher contrastive loss. Thus, minimizing Eq. (7) will drive the model to update towards making more accurate mask predictions, i.e., rectifying the bias of CLIP when adapting CLIP to the downstream USS task.
Note that in ReCLIP, an additional distillation stage is employed, where the knowledge of rectified CLIP is distilled into a specifically designed segmentation network like DeepLab-v2 [68]. However, in ReCLIP++, we don’t need such a distillation stage. Empirically, we observe that such a distillation process has two opposing effects. On the one hand, it may benefit the smoothness of the predicated mask, which has a positive effect on the segmentation performance. On the other hand, instead of mitigating the bias, it may enlarge the bias to an extent, which has a negative effect on the segmentation performance. Benefitting from $\mathcal{F}_{dec}$, the rectified mask of ReCLIP++ is much smoother and contextual than that in ReCLIP, rendering the negative effect of distillation outweigh the positive effect of distillation. Thus, we choose not to employ the distillation in ReCLIP++ to make the training more simplified and achieve even better segmentation performance (see Sec. 4.4 for more details).
3.5 Inference
@@
For the inference with ReCLIP++, we obtain the query logit map $M_{q}$ by the query segmentation head and the bias logit map $M_{b}$ which models both biases by the Bias Extraction Module. We then obtain the rectified output $M_{o}$ by the Rectified Mask Generation Module with Eq. (4). With a simple $\arg\max$ operation to $M_{o}$ across the class channel $C$, we generate the rectified masks as final predictions.
4 Experiment
4.1 Setup
@@
Datasets. We conduct experiments on five standard benchmarks for semantic segmentation, including PASCAL VOC 2012 [25], PASCAL Context [26], ADE20K [27], Cityscapes [28] and COCO Stuff [29]. PASCAL VOC 2012 (1,464/1,449 train/validation) contains 20 object classes, while PASCAL Context (4,998/5,105 train/validation) is an extension of PASCAL VOC 2010 and we consider 59 most common classes in our experiments. ADE20K (20,210/2,000 train/validation) is a segmentation dataset with various scenes and 150 most common categories are considered. Cityscapes (2,975/500 train/validation) consists of various urban scene images of 19 categories from 50 different cities. COCO Stuff (118,287/5,000 train/validation) has 171 low-level thing and stuff categories excluding background class. Following the previous evaluation protocol of unsupervised semantic segmentation methods [21, 46], we use 27 mid-level categories for training and inference.
@@
Implementation details. For the image encoder of CLIP [19], we adopt ViT-B/16 as our visual backbone. For the text encoder of CLIP, we adopt Transformer [69]. During the whole training period, we keep both of the encoders frozen. We use conventional data augmentations including random cropping and random flipping. We use an SGD [70] optimizer with a learning rate of 0.01 and a weight decay of 0.0005. We adopt the poly strategy with the power of 0.9 as the learning rate schedule. We set $r=1/6$ for all the datasets. We set $t=7\%$ for PASCAL VOC and Cityscapes, and $t=0$ for the rest datasets. In our experiment, we report the mean intersection over union (mIoU) as the evaluation metric. More details about training can be found in our appendix.
4.2 Comparison with previous SOTA